# Rollbaccine: Herd Immunity against Storage Rollback Attacks in TEEs (Technical Report)

### v1

@@ Abstract.
Today, users can “lift-and-shift” unmodified applications into modern, VM-based Trusted Execution Environments (TEEs) in order to gain hardware-based security guarantees. However, TEEs do not protect applications against disk rollback attacks, where persistent storage can be reverted to an earlier state after a crash; existing rollback resistance solutions either only support a subset of applications or require code modification. Our key insight is that restoring disk consistency after a rollback attack guarantees rollback resistance for any application. We present Rollbaccine, a device mapper that provides automatic rollback resistance for all applications by provably preserving disk consistency. Rollbaccine intercepts and replicates writes to disk, restores lost state from backups during recovery, and minimizes overheads by taking advantage of the weak, multi-threaded semantics of disk operations. Across benchmarks over two real applications (PostgreSQL and HDFS) and two file systems (ext4 and xfs), Rollbaccine adds only $19\%$ overhead, except for the fsync-heavy Filebench Varmail. In addition, Rollbaccine outperforms the state-of-the-art, non-automatic rollback resistant solution by 208×208$\times$208 ×.
copyright: cc
@@ 1. Introduction
Security-conscious developers lift-and-shift unmodified applications into VM-based Trusted Execution Environments (TEEs) under the impression that TEEs guarantee confidentiality and integrity with minimal performance overhead (SNPPerfPost, ; SNPPerfPost2, ; cockroachCVM, ).
@@
This is true until the application needs to access disk; TEEs only protect data in memory, leaving the disk vulnerable. A combination of encryption, sealing, and hash verification can be used to provide confidentiality and integrity while the host is online, but once the host goes offline, the data on disk become vulnerable to rollback attacks.
@@
Rollback attacks revert disk to an earlier state, causing the system to execute over stale data. Such attacks can be devastating: for example, an attacker can use rollback attacks in order to bypass limits on password attempts (androidRollbackKeys, ; iosRollbackKeys, ; signalRollbackKeys, ; juiceboxRollbackKeys, ) or reopen vulnerabilities in patched software (codeTransparencyService, ; windowsDowngradeAttack, ). To combat rollback attacks, existing TEE-based systems in production must abandon the lift-and-shift philosophy in order to implement bespoke rollback protection mechanisms (svr3, ; ccf, ; gcpSpaces, ; azureConfidentialComputing, ; awsDmVerity, ).
Ideally, there would exist a solution against rollback attacks that is at once (1) general, correct for all applications, (2) automatic, requiring no application modification, and (3) resistant, allowing the application to recover as if the rollback attack did not occur. Unfortunately, despite the variety of existing solutions against rollback attacks (rote, ; nimble, ; narrator, ; treaty, ; speicher, ; secureFS, ; engraft, ; svr3, ; vpfs, ; mlsdisk, ; dm-x, ; restartRollback, ; memoir, ; ariadne, ; ccf, ), none achieve all three properties. Existing solutions either only detect but do not recover from rollbacks (treaty, ; speicher, ; secureFS, ; vpfs, ; mlsdisk, ; dm-x, ; memoir, ; ariadne, ), are application specific (ccf, ; engraft, ; svr3, ; narrator, ), or sacrifice automation by requiring the application to use a new API (nimble, ).
@@
In this paper, we make one key observation: rollback attacks are fundamentally attacks on disks. Therefore, generality can be achieved by restoring disk consistency after rollback, guaranteeing rollback resistance to any application that uses disk regardless of application semantics.
@@
The key challenge then lies in developing a strategy that preserves disk consistency at low cost. Replication will necessarily be part of the solution: at least one machine must still have the data! Naïvely replicating all disk updates during execution, however, is a non-starter performance wise; this is why Nimble, the state-of-the-art solution, requires the application to use a new API to indicate when replication is necessary (nimble, ).
@@
A new API is unnecessary. Already encoded in the semantics of disk operations are persistence flags (REQ_FUA and REQ_PREFLUSH(fuaPreflush, ), metadata attached to each write request indicating whether it should be synchronously written to disk or not. Disk writes without these flags can return before persistence is guaranteed and potentially be lost after a crash. These semantics are already used by file systems in order to make most writes to disk fast and asynchronous by default, while a small number of operations are carefully persisted to ensure correctness. We can build off of these semantics when replicating disk for rollback resistance; writes with persistence flags must be replicated on the critical path, while all other writes can be replicated in the background.
@@
The weak semantics of disk also allow us to relax constraints on ordering. All existing countermeasures against rollback attacks enforce a total ordering of state changes in order to identify a “canonical” state that the system must recover to. Disks are more flexible. Upon crash and recovery, disks can recover any subset of weakly-persisted writes. We can take advantage of this flexibility when replicating disk, allowing each disk to process writes in different orders and diverge, as long as they remain in a state consistent with prior operations.
@@
We instantiate these ideas in Rollbaccine (the rollback vaccine), a system that intercepts and replicates writes to provide general and automatic rollback resistance with minimal overhead. To prove that Rollbaccine restores disk consistency, we formally define the behavior of block devices (a category of storage devices that includes disk) in the presence of crashes (§ 4) and prove that block device consistency is preserved by Rollbaccine (Appendix B).
@@
Our experimental results confirm that with Rollbaccine, general and automatic rollback resistance is possible with minimal performance penalty. On two large applications, PostgreSQL and HDFS, as well as two distinct file systems, ext4 and xfs, Rollbaccine introduces a maximum of $19\%$ throughput and latency overhead across TPC-C (tpc-c, ), NNThroughputBenchmark (nnThroughputBenchmark, ), and Filebench (filebench, ) Webserver, with more significant overheads ($71\%$ throughput and 2.7×2.7$\times$2.7 × latency) only for the Filebench Varmail benchmark, with its high fsync frequency. In addition, Rollbaccine outperforms Nimble (nimble, )—a state-of-the-art, non-automatic rollback resistance solution—by 208×208$\times$208 × for write-heavy workloads.
@@
    We show that Rollbaccine adds minimal overhead in most benchmarks and significantly outperforms state-of-the-art, non-automatic rollback-resistant solutions (§ 7).
@@ 2. Motivation and Threat Model
TEEs protect the confidentiality and integrity of applications by preventing, through hardware and software, unauthorized access to code or data. Users can verify that their applications are executing within a TEE through remote attestation, where the host produces a proof of the code executing within a TEE (attestation, ).
Until recently, applications that wished to run within TEEs (e.g. Intel SGX) required extensive modifications and suffered significant performance penalties (scone, ). VM-based TEEs such as Intel TDX (tdx, ), AMD SEV-SNP (SEV-SNP, ), and Arm CCA (ARM-CCA, ) provide a new “lift-and-shift” abstraction, where applications can run unmodified inside the TEE and gain confidentiality and integrity guarantees with minimal performance overhead. TEEs have seen widespread adoption as a result: all major cloud providers support at least one type of VM-based TEE (azureConfidentialComputing, ; googleConfidentialComputing, ; awsConfidentialComputing, ), and they are used in industry for private data processing (gcpSpaces, ; ccr, ; edgeless-mpc, ), key management (signalRollbackKeys, ), supply-chain security (codeTransparencyService, ), and AI inference (azureConfidentialAI, ; nvidiaEdgeless, ; federatedLearning, ).
@@ 2.1. The dangers of rollbacks
Unfortunately, the confidentiality and integrity guarantees do not currently extend to persistent state. Attackers can observe and modify application data on disk by either directly accessing disk (using other applications) or by intercepting disk operations with privileged code (such as a kernel module loaded into the host OS).
Existing encryption and integrity-preserving techniques (dmCrypt, ; dmIntegrity, ) can be used to automatically provide disk confidentiality and some degree of integrity. However, because the metadata used to verify integrity is still stored on disk, they remain vulnerable to rollback attacks, where an adversary could modify data (and its on-disk integrity metadata, in tandem) in order to present the application with a stale disk.
@@
Online rollback attacks—performed while the application is executing—can be detected by an application that validates reads against integrity metadata in memory. An attacker can instead launch an offline rollback attack, crashing the TEE (and thereby clearing any metadata in memory) before rolling disk back. Offline rollback attacks are insidious because they are undetectable; the recovering application cannot distinguish an offline rollback attack from a benign crash, and will execute obliviously on stale state.
Consider for example a TEE application that rate-limits password-guessing attempts (LABEL:lst:password-guessing). It maintains a counter on disk to prevent excessive retries, even across reboots. An attacker could repeatedly crash the TEE, rollback the disk to a state before the counter was incremented, then restart the TEE, effectively offering users unlimited password-guessing attempts.
@@
if counter > 10:
print("Accountblocked")
@@
Rollback attacks are relevant to all applications that rely on persistent storage, including applications that do not interface with disk directly and instead rely on local (or even distributed!) database systems for persistence. Those systems in turn rely on the persistence of their local disks, which can be violated by rollback attacks.
To combat rollback attacks, production-level TEE-based systems all implement bespoke rollback-detecting or resistant solutions. SVR3 (svr3, ), Signal’s key recovery system, modifies Raft (raft, ) to prevent rollback attacks. CCF (ccf, ), Azure’s TEE-protected ledger, constructs a Merkle tree to prevent rollbacks. Google’s Confidential Space (gcpSpaces, ), Azure’s Confidential Containers (azureConfidentialComputing, ), and AWS’s Bottlerocket (awsDmVerity, ) all verify the integrity of disk to detect modifications.
@@ 2.2. Threat model and guarantees
Threat model. We make the standard assumptions, common to all TEE-based systems (nimble, ; ccf, ; restartRollback, ; narrator, ; serverlessCC, ; cyclosa, ), that clients trust the hardware manufacturer and the TEE, and that TEEs are as safe as they claim to be. Specifically, attackers cannot violate the integrity and confidentiality of memory, break standard cryptographic primitives, or exploit physical hardware or side-channel attacks (Gruss17, ; Oleksenko18, ; Van20, ; Van18, ; SGAxe, ; Borrello22, ; Murdock19, ; Skarlatos19, ; Schwarz17, ; VanSchaik22, ; asyncShock, ). Attackers, including malicious cloud providers, can still crash machines and corrupt network and disk I/O.
@@
An application is rollback-detecting if it always detects rollback attacks.
Rollback detection provides safety. Following a rollback attack, a rollback-detecting application may have lost its data and be incapable of recovering, but it will never execute over stale data. In other words, rollback detection transforms rollback attacks into denial-of-service attacks.
@@
Rollback resistance provides liveness in addition to safety. Following a rollback attack, a rollback-resistant application will recover enough lost data to continue execution.
Non-guarantees. Rollback resistance guarantees that the effect of a rollback is invisible to the application. It does not, however, guarantee correctness.
Consider again the password guessing application. If the application did not include the os.fsync line, an attacker could intelligently crash the application before the counter makes it to disk, thus bypassing the password guessing limit without ever modifying disk. Because the attacker did not rollback the disk, this is not a rollback attack and remains possible even if the application were rollback resistant. Similarly, rollback resistance’s guarantees are limited to disk operations; it does not provide general safety to arbitrary applications, e.g. if the password counter relies on accurate system time from the untrusted host, or uses an insecure entropy source for password generation.
@@ 3. Towards Rollbaccine
Generality—the ability to protect arbitrary applications against rollback attacks—is the hardest to achieve, because we do not know what data each application relies on for recovery. To achieve generality, solutions like Nimble (nimble, ) sacrifice automation, requiring significant manual effort to identify the application-specific state that must be protected.
@@
Rollbaccine leverages this insight to implement general, automatic rollback resistance by replicating disk. Replication is a well-known strategy for recovering data lost in a rollback attack (nimble, ; narrator, ; rote, ; engraft, ). Replicating disk, however, requires addressing three main challenges: (1) understanding exactly what state the disk can recover to after a benign crash, (2) intercepting write operations to disk and replicating them so they are available after a rollback attack, and (3) limiting the overheads of doing so. In the rest of this paper, we address each challenge in turn:
1. Formalisms (§ 4). In order to restore disk to a state it could have recovered to after a benign crash, we need a formal understanding of exactly what states it could have possibly been in. To the best of our knowledge, although file system crash consistency (ace, ; crashmonkey, ; alice, ; crashMonkeyAndACE, ; optimisticCrashConsistency, ; vijayThesis, ; daisyNFS, ; hydra, ; chipmunk, ; yggdrasil, ; ferrite, ; monarch, ; vijayThesis, ) has been extensively studied, there is no formal definition for the block-level semantics of disks.
f = open(”sosp.txt”) WRITE(8)
@@
Table 1. Opening, writing, and fsyncing to a file in ext4 and their corresponding block I/Os, sector numbers, and persistence flags.
2. Intercepting (§ 6). Next, we present the implementation of Rollbaccine as a device-mapper. A Linux device mapper is a kernel module that lies between block devices (such as disks) and higher level applications, as seen in Figure 1. Each disk read or write request arrives at the device mapper as a block I/O consisting of the disk sectors involved, pages containing data for writes or retrieving data for reads, and additional flags (REQ_FUA and REQ_PREFLUSH) describing whether the data should be written synchronously to disks or not. Table 1 describes a simple application writing to the sosp.txt file on the ext4 file system and the resulting block I/Os.
Implementing Rollbaccine as a device mapper presents three benefits. First, it is a commonly accepted strategy in industry to augment disk functionality. Dm-crypt (dmCrypt, ), dm-verity (dmVerity, ), and dm-integrity (dmIntegrity, ) are all popular device mappers industry used to enable disk encryption and (limited) integrity without application modification (awsDmVerity, ; azureDmCrypt, ; azureDmVerity, ; azureStorage, ; androidDmCrypt, ; androidDmVerity, ). Second, device mappers sit below the file system (Figure 1) and are thus file-system agnostic; this allows us to evaluate against both ext4 and xfs in § 7.1 without code modification. Third, device mappers have a well defined interface that is relatively simple to review, maintain, optimize, and trust, as opposed to a custom rollback-resistant port of an existing application (nimble, ) or a custom rollback resistant file system.
3. Overheads (§ 6). Finally, we discuss how Rollbaccine minimizes performance overheads. Synchronous disk replication on the critical path not only introduces high overhead (§ 7.2), but is also often unnecessary. Existing applications and file systems already carefully engineer their implementation to reduce the number of synchronous writes; these writes are mapped to block I/Os with persistence flags and exposed to the device mapper. It suffices for Rollbaccine to synchronously replicate writes with persistence flags and replicate remaining writes in the background.

### v2

@@ Abstract.
Today, users can “lift-and-shift” unmodified applications into modern, VM-based Trusted Execution Environments (TEEs) in order to gain hardware-based security guarantees. However, TEEs do not protect applications against disk rollback attacks, where persistent storage can be reverted to an earlier state after a crash; existing rollback resistance solutions either only support a subset of applications or require code modification. Our key insight is that restoring disk consistency after a rollback attack guarantees rollback resistance for any application. We present Rollbaccine, a device mapper that provides automatic rollback resistance for all applications by provably preserving disk consistency. Rollbaccine intercepts and replicates writes to disk, restores lost state from backups during recovery, and minimizes overheads by taking advantage of the weak, multi-threaded semantics of disk operations. Rollbaccine performs on-par with state-of-the-art, non-automatic rollback resistant solutions; in fact, across benchmarks over PostgreSQL, HDFS, and two file systems (ext4 and xfs), Rollbaccine adds only $19\%$ overhead, except for the fsync-heavy Filebench Varmail.
copyright: $c$
@@ 1. Introduction
Security-conscious developers lift-and-shift unmodified applications into VM-based Trusted Execution Environments (TEEs) under the impression that TEEs guarantee confidentiality and integrity with minimal performance overhead (SNPPerfPost; SNPPerfPost2; cockroachCVM) [^1] It’s important to highlight that TEEs only protect confidentiality and integrity. An insecure application due to a code-level bug remains insecure.. This is true until the application needs to access disk; TEEs only protect data in memory, leaving the disk vulnerable. A combination of encryption, sealing, and hash verification can be used to provide confidentiality and integrity while the host is online, but once the host goes offline, the data on disk becomes vulnerable to rollback attacks.
@@
Rollback attacks revert disk to an earlier state, causing the system to execute over stale data. Such attacks can be devastating: for example, an attacker can use rollback attacks in order to bypass limits on password attempts (androidRollbackKeys; iosRollbackKeys; signalRollbackKeys; juiceboxRollbackKeys) or reopen vulnerabilities in patched software (codeTransparencyService; windowsDowngradeAttack).
To combat rollback attacks, production-level TEE-based systems implement bespoke rollback-detecting or resistant solutions. Signal’s SVR3 (svr3) modifies Raft (raft) while Azure’s CCF (ccf) constructs a Merkle tree, whereas Google’s Confidential Space (gcpSpaces), Azure’s Confidential Containers (azureConfidentialComputing), and AWS’s Bottlerocket (awsDmVerity) verify code integrity at launch. Academic efforts have further extended the catalog of supported applications (rote; nimble; narrator; treaty; speicher; secureFS; engraft; vpfs; mlsdisk; dm-x; restartRollback; memoir; ariadne).
Lift-and-shift rollback resistance, however, remains elusive. There is no solution that is at once (1) general, correct for any application, (2) automatic, requiring no application modification, and (3) rollback resistant, allowing the application to recover as if the rollback attack did not occur.
@@
We find in this paper that general, automatic rollback resistance is not only possible, but that its performance is also competitive against state-of-the-art, manual solutions. Our key observation is that rollback attacks are fundamentally attacks on disk consistency. General rollback resistance can therefore be achieved by restoring disk consistency after rollback, guaranteeing rollback resistance to any application that uses disk regardless of application semantics.
@@
The key challenge then lies in developing a strategy that preserves disk consistency at low cost. Replication will necessarily be part of the solution: at least one machine must still have the data! Naïvely replicating all disk updates during execution, however, is a non-starter performance wise; this is why Nimble (nimble), the state-of-the-art solution, requires the application to use a new API to indicate when replication is necessary.
@@
A new API is unnecessary; this information is already available at the Linux block device layer (Figure 1). Disk operations already include persistence flags (REQ_FUA and REQ_PREFLUSH(fuaPreflush), metadata attached to each write request indicating whether it should be synchronously written to disk or not. Disk writes without these flags can return before persistence is guaranteed and potentially be lost after a crash. These semantics are already used by file systems in order to make most writes to disk fast and asynchronous by default, while a small number of operations are carefully persisted to ensure correctness. We can build off these semantics when replicating disk for rollback resistance; writes with persistence flags must be replicated on the critical path, while all other writes can be replicated in the background.
@@
The weak semantics of disk also allow us to relax constraints on ordering, further improving performance. All existing countermeasures against rollback attacks enforce a total ordering of state changes in order to identify a “canonical” state that the system must recover to. Disks are more flexible. Upon crash and recovery, disks can recover any subset of weakly-persisted writes. We can take advantage of this flexibility when replicating disk, allowing each disk to process writes in different orders and diverge, as long as they remain in a state consistent with prior operations.
@@
We instantiate these ideas in Rollbaccine (the rollback vaccine), a system that intercepts and replicates writes to provide general and automatic rollback resistance with minimal overhead. To prove that Rollbaccine restores disk consistency, we formally define the behavior of block devices (a category of storage devices that includes disk) in the presence of crashes and prove that block device consistency is preserved by Rollbaccine.
@@
Our experimental results confirm that with Rollbaccine, general and automatic rollback resistance is possible with minimal performance penalty. Across applications (PostgreSQL, HDFS, and the most common Linux file systems ext4 and xfs) and benchmarks (TPC-C (tpc-c), NNThroughputBenchmark (nnThroughputBenchmark), Filebench (filebench)) Rollbaccine introduces a maximum of $19\%$ throughput and latency overhead (on-par with Nimble (nimble), a state-of-the-art, non-automatic rollback resistance solution) with more significant overheads ($71\%$ throughput and $2.7\times$ latency) only for Filebench Varmail, with its high fsync frequency.
@@
    We show that Rollbaccine adds minimal overhead in most benchmarks and is comparable to state-of-the-art, non-automatic rollback-resistant solutions (§ 7).
@@ 2. Motivation and Threat Model
TEEs provide confidentiality and integrity guarantees by preventing, through hardware and software, unauthorized access to code or data. Users can verify that their applications are executing within a TEE through remote attestation, where the host produces a proof of the code executing within a TEE (attestation).Until recently, applications that wished to run within TEEs (e.g. Intel SGX) required extensive modifications and suffered significant performance penalties (scone).
VM-based TEEs such as Intel TDX (tdx), AMD SEV-SNP (SEV-SNP), and Arm CCA (ARM-CCA) provide a new “lift-and-shift” abstraction, where applications can run unmodified inside the TEE with minimal performance overhead. TEEs have seen widespread adoption as a result: all major cloud providers support at least one type of VM-based TEE (azureConfidentialComputing; googleConfidentialComputing; awsConfidentialComputing), and they are used in industry for private data processing (gcpSpaces; ccr; edgeless-mpc), key management (signalRollbackKeys), supply-chain security (codeTransparencyService), and AI inference (azureConfidentialAI; nvidiaEdgeless; federatedLearning).
@@ 2.1. The Dangers of Rollbacks
Unfortunately, the confidentiality and integrity guarantees do not currently extend to persistent state. Existing encryption and integrity-preserving techniques (dmCrypt; dmIntegrity; bitlocker) can be used to automatically provide disk confidentiality and some degree of integrity. However, because the metadata used to verify integrity is still stored on disk, they remain vulnerable to rollback attacks, where an adversary could modify data (and its on-disk integrity metadata, in tandem) in order to present the application with a stale disk.
@@
Online rollback attacks—performed while the application is executing—can be detected by an application that validates reads against integrity metadata in memory. An attacker can instead launch an offline rollback attack, crashing the host (and thereby clearing any metadata in memory) before rolling disk back. Offline rollback attacks are insidious because they are undetectable; the recovering application cannot distinguish an offline rollback attack from a benign crash, and will execute obliviously on stale state.
Consider for example a TEE application that rate-limits password-guessing attempts (LABEL:lst:password-guessing). It maintains a counter on disk to prevent excessive retries, even across reboots. An attacker could repeatedly crash the TEE and rollback the disk to a state before the counter was incremented, effectively bypassing the guessing limit.
@@
if counter > 10: # Too many attempts
@@
Rollback attacks are relevant to all applications that rely on persistent storage, including applications that do not interface with disk directly and instead rely on local (or even distributed!) database systems for persistence. Those databases in turn rely on the persistence of their local disks, which can be violated by rollback attacks.
@@ 2.2. Threat Model and Guarantees
Threat model. We make the standard assumptions that clients trust the hardware manufacturer, and that TEEs are as safe as they claim to be (nimble; ccf; restartRollback; narrator; serverlessCC; cyclosa). Specifically, attackers cannot violate the integrity and confidentiality of memory, break standard cryptographic primitives, or exploit physical hardware or side-channel attacks (Gruss17; Oleksenko18; Van20; Van18; SGAxe; Borrello22; Murdock19; Skarlatos19; Schwarz17; VanSchaik22; asyncShock). Applications executed within a TEE will not deviate from their code. Attackers can still crash machines and corrupt network and disk I/O.
As in any system, the number of machines that an attacker can compromise directly impacts the correctness guarantees that the system can make; we therefore classify attackers as either Type I (Individual) or Type II (Total). Type I attackers represent malicious employees that can compromise up to $f$ machines, while Type II attackers represent malicious cloud providers with the ability to compromise all machines.
@@
An application is rollback-detecting if it detects rollback attacks and halts.
Rollback detection guarantees safety. Following a rollback attack, a rollback-detecting application may be incapable of recovering, but it will never execute over stale data.
@@
Rollback resistance guarantees liveness in addition to safety. Following a rollback attack, a rollback-resistant application will recover lost data and continue execution.
Rollbaccine guarantees rollback resistance in the presence of Type I attackers and rollback detection for Type II attackers.
2.3. Limitations
Type II attackers can always violate liveness (and rollback resistance) by simply denying service or shutting down all machines; this is a fundamental limitation of all rollback solutions  (nimble; secureKeeper; lcm; cyclosa). In that case, Rollbaccine falls back to rollback detection, ensuring that the application does not execute over tampered data. For clients wary of cloud providers, a multi-cloud deployment would effectively convert all Type II attackers to Type I.
Importantly, rollback resistance should not be equated with application correctness. An application with preexisting vulnerabilities, placed in a TEE and given rollback resistance, does not become secure. Probabilistically unlikely failure modes in an application (or the underlying file system) may be exploited through non-rollback attacks.
Consider again the password-guessing application with the os.fsync line removed. It is probabilistically correct in a benign setting where crashes are rare; however, an attacker could repeatedly crash the TEE before the counter makes it to disk in order to bypass the password guessing limit. This is not a rollback attack—the attacker crashed the TEE but did not modify the disk—and remains possible even with rollback resistance. Rollback resistance simply removes the effect of rollback attacks.
@@ 3. Towards Rollbaccine
Generality—the ability to protect arbitrary applications against rollback attacks—is the hardest to achieve, because we do not know what data each application relies on for recovery. To achieve generality, solutions like Nimble (nimble) sacrifice automation, requiring significant manual effort to identify the application-specific state that must be protected.
@@
Rollbaccine leverages this insight to implement general, automatic rollback resistance by replicating disk. Replication is a well-known strategy for recovering data lost in a rollback attack (nimble; narrator; rote; engraft). Replicating disk, however, requires addressing three main challenges: (1) understanding exactly what states the disk can recover to after a benign crash, (2) intercepting and replicating writes to disk so they are available after a rollback attack, and (3) limiting the overheads of doing so. In the rest of this paper, we address each challenge in turn:
1. Formalisms (§ 4). In order to restore disk to a state it could have recovered to after a benign crash, we need a formal understanding of exactly what states it could have possibly been in. Although file system crash consistency has been extensively studied (ace; crashmonkey; alice; crashMonkeyAndACE; optimisticCrashConsistency; daisyNFS; hydra; chipmunk; yggdrasil; ferrite; monarch; vijayThesis), the field is primarily concerned with avoiding inconsistent states or detecting violations through testing. To the best of our knowledge, no existing work formalizes the correct behaviors for the disk that these file systems rely on.
f = open(”test.txt”) WRITE(8)
@@
Table 1. File operations in ext4 and their corresponding block I/Os, sector numbers, and persistence flags.
2. Intercepting (§ 6). Next, we need a well defined interface for intercepting disk operations that is relatively simple to review, maintain, optimize, and trust, as opposed to a custom rollback-resistant application (nimble) or file system. To this end, we implement Rollbaccine as a device-mapper.
A Linux device mapper is a kernel module that lies between block devices (such as disks) and higher level applications, as seen in Figure 1. Each disk read or write request arrives at the device mapper as a block I/O consisting of the disk sectors involved, pages containing data for writes or retrieving data for reads, and additional flags (REQ_FUA and REQ_PREFLUSH) describing whether the data should be written synchronously to disks or not. Table 1 describes a simple application writing to the test.txt file on the ext4 file system and the resulting block I/Os.
Implementing Rollbaccine as a device mapper presents two benefits. First, it is a commonly accepted strategy in industry to augment disk functionality. Dm-crypt (dmCrypt), dm-verity (dmVerity), and dm-integrity (dmIntegrity) are all popular device mappers for enabling disk encryption and (limited) integrity without application modification (awsDmVerity; azureDmCrypt; azureDmVerity; azureStorage; androidDmCrypt; androidDmVerity). Second, device mappers sit below the file system (Figure 1) and are thus file-system agnostic; this allows us to evaluate against both ext4 and xfs in § 7.1 without code modification.
3. Overheads (§ 6). Finally, we must minimize the overhead of replication. Synchronous disk replication on the critical path not only introduces high overhead (§ 7.2), but is also often unnecessary. Existing applications and file systems already carefully engineer their implementation to reduce the number of synchronous writes; these writes are mapped to block I/Os with persistence flags and exposed to the device mapper. It suffices for Rollbaccine to synchronously replicate writes with persistence flags and replicate remaining writes in the background.